OpenStudy (anonymous):

How do you solve 8x^3-27?

OpenStudy (precal):

it needs to be equal to something for you to solve it

OpenStudy (anonymous):

its equal to zero, sorry

OpenStudy (lgbasallote):

convert it to the form (a^3 + b^3) = (a+b)(a^2-ab + b^2)

OpenStudy (lgbasallote):

(2x + 3)(4x^2 - 6x + 9)...just distribute to get the final answer

OpenStudy (anonymous):

so do i set the first equation to zero and solve then use the quadratic formula to solve the second equation to find the roots?

OpenStudy (lgbasallote):

no! distribute....

OpenStudy (anonymous):

i'm lost

OpenStudy (lgbasallote):

oh wait...i was factoring :))) I should've been solving. Sorry for confusing you. Transpose 27 to the other side. 8x^3 = 27 get the cube root 2x = 3 x = 3/2

OpenStudy (anonymous):

there should be two more roots..

OpenStudy (anonymous):

that's okay, thank-you

OpenStudy (lgbasallote):

@cinar...i think this is a repetitive root

OpenStudy (anonymous):

no, they are complex

OpenStudy (anonymous):

\[\frac34(-1+i \sqrt{3})\quad and \quad \frac34(-1-i \sqrt{3})\quad are \,also\, roots\]